Calculate the number of atom of Oxygen Present in 8 gram of Co2.what would the weight of CO having the same number of oxygen atoms

Grade:11

1 Answers

Vikas TU
7 years ago
First calculate the moles of CO2.
moles = 8/44 = 2/22 = 1/11 mol.
thus, moles of O atom = 2/11
No. of atoms = 2/11 * Na
Where Na is avogadro number.